OpenStudy (anonymous):

Which value of x makes the following equation true? 2(x - 2) = -3x + 26 A. 7 B. 6 C. 9 D. 10

OpenStudy (anonymous):

So try the same thing we did before starting with x = 7

OpenStudy (campbell_st):

just substitute and check the left hand side is equal to the right hand side

OpenStudy (anonymous):

2(7-2) = -3(7) + 26 Then do x = 6 2(6-2) = -3(6) + 26 etc! Until one side is the same as the other.

OpenStudy (anonymous):

So B?